&. The pie chart below shows monthly the budget for a Family. Other Epenses 8% Grocery 2196 Saving 23% Automobile 22% Healthcare 6% Dining Out 8% / uailitv Bills 12% What percentage of the income does the family spend on healthcare and utility bills? a. $29 b. $360 c. $580 d. $420
Using the table below to answer questions 3 and 4 The table shows the results of a survey of students in two STAT 101 classes. Did You Watch More Than One Hour of TV Last Night? Has 8 am classYes No Yes 0 13 10 3 What is the probability that a student watch more than one hour TV or has no 8 am class? a. 0.85 4 Find P(more than 1 hour of TV | Has 8am class) a. 0.647 5. A student claims that the averan b. 0 c. 0.25 d. 0.75 b. 0.565 с. 0.435 d. 0.765
Answer #1

Ans 3

P= favourable case/total number of cases

P = (11+13+10)/40

a. 0.85

Ans 4

P = 11/17 =

a. 0.647